# Mastering Collaborative Multi-modal Data Selection: A Focus on Informativeness, Uniqueness, and Representativeness

### v1

@@ Abstract
Instruction tuning fine-tunes pre-trained Multi-modal Large Language Models (MLLMs) to handle real-world tasks. However, the rapid expansion of visual instruction datasets introduces data redundancy, leading to excessive computational costs. We propose a collaborative framework, DataTailor, which leverages three key principles—–informativeness, uniqueness, and representativeness–—for effective data selection. We argue that a valuable sample should be informative of the task, non-redundant, and represent the sample distribution (i.e., not an outlier). We further propose practical ways to score against each principle, which automatically adapts to a given dataset without tedious hyperparameter tuning. Comprehensive experiments on various benchmarks demonstrate that DataTailor achieves 100.8% of the performance of full-data fine-tuning with only 15% of the data, significantly reducing computational costs while maintaining superior results. This exemplifies the “Less is More" philosophy in MLLM development.
@@ 1 Introduction
The rapid development of Multi-modal Large Language Models (MLLMs) has made promising progress on various multi-modal tasks [61, 3, 14, 26, 67]. A typical MLLM is developed through two main training stages: pre-training on vast image-text pairs and fine-tuning on task-specific multi-modal instructions. Notably, the fine-tuning stage is critical for enhancing the instruction-following capabilities of MLLMs. Yet, this stage can become exceedingly time-consuming due to the large-scale but low-quality instruction data. Hence the community is interested in fine-tuning data selection methods, such that an MLLM trained on the selected subset yields comparable or even better performance.
Figure 1: (a) The Performance v.s. Selected Data Ratio on LLaVA-mix-665k of DataTailor compared with SOTA data selection methods. (b) Metric triangle among informativeness, uniqueness, and representativeness when applying IFD [30] (heuristic score methods), InsTag [37] (human-reward methods), LESS [54] (gradient-based methods), and our DataTailor for multi-modal data selection. (c) The corresponding MLLM performance on LLaVA-mix-665k [34] of different data selection methods.
Existing MLLM data selection methods [10, 53, 36, 19] largely follow similar ideas from the NLP community [30, 37, 66, 47, 54]. They can be divided into three main categories: (1) Heuristic score methods [30, 11, 7] leverage pre-defined rules to select data, which are not flexible to handle diverse downstream tasks. (2) Human-reward methods [66, 37] utilize human feedback to select data, which are both time-consuming and expensive. (3) Gradient-based methods [47, 54, 32] select samples whose training loss gradient aligns with that averaged over the dataset. However, they are unable to filter redundant samples.
To address these deficiencies, we build a systematic data selection method for MLLM called DataTailor. We evaluate each sample with three principles and select the most valuable samples, leading to state-of-the-art MLLM performance with a fraction of data (c.f. Fig. 1). The three principles are: (1) Informativeness: a valuable sample should be informative of the task, e.g., If the task is reasoning, describing the differences in movements between skiing and ice skating is more informative than simply showing someone skiing. In Fig. 2 where each axis (heuristically) represents an orthogonal dimension of task information, points along the diagonal carry more information about the task. (2) Uniqueness: a valuable sample should be distinct from others, exhibiting a large distance from nearby samples to reduce data redundancy (c.f. Fig. 2 near the blue dashed region in the intra-cluster space demonstrate high uniqueness). (3) Representativeness: it should be a typical sample in the data distribution. This prevents selecting noisy outliers or mislabeled samples (c.f. Fig. 2 the clusters connected by blue lines in the inter-cluster space exhibit high representativeness for the overall dataset).
@@
We further propose a practical method to measure the value of each sample against each principle. For informativeness, we take motivations from information theory [9, 8]. For each sample, we analyze the singular value distribution of its features and use the entropy of the singular values to determine if it is informative of the task. To compute uniqueness and representativeness, we first cluster the samples based on their visual and textual features. This allows efficient calculation for uniqueness, as we can simply measure the average distances of a sample to its neighbors in the same cluster, and mark those with a large distance as unique. Then we find connected clusters and mark samples in those clusters as representative. Hence noisy or mislabeled data from a far-away cluster can be filtered.
Moreover, as multi-modal samples exhibit varying structures and complexities across diverse tasks, we propose an adaptive weight to combine the values, which removes the need for expensive hyper-parameter tuning. We also adaptively determine the proportion of selected data for each task by using the average largest singular value of samples, which empirically reflects task difficulty and correlates with training convergence. Combining these techniques, DataTailor synergizes the three principles for data selection and achieves an optimal balance between data volume and model performance (as shown Fig. 1(a) red line).
To our knowledge, we are the first to explore sample relationships between multi-modal instructions systematically. Through extensive experiments, we demonstrate that DataTailor exhibits significant effectiveness in data selection for MLLMs (with less than 5% data but achieving over 95% performance). This effectiveness stems from our comprehensive evaluation based on the three core principles, ensuring that the selected data excels in all three aspects (c.f. Fig. 1(b)). In contrast, other methods lack a systematic evaluation, particularly of sample relationships, leading to weaknesses in uniqueness and representativeness and resulting in suboptimal MLLM performance (c.f. Fig. 1(c)). Remarkably, when DataTailor increases the data selection ratio, multi-modal data selection can even outperform full data fine-tuning (achieve 100.8% performance with 15% data), truly exemplifying the concept of “Less is More". Overall, our main contributions are summarized as follows:
@@ 2.1 Multi-modal Large Language Model
With the outstanding performance of LLMs in zero-shot settings, early work combining LLMs with visual modalities has demonstrated impressive visual language comprehension abilities [15, 61, 28, 25, 29, 39, 49, 40, 62, 23, 27]. Recently, more powerful MLLMs have emerged [67, 41, 60, 10, 34, 6, 59, 17], which possess perceptual abilities for visual-language tasks and excellent reasoning abilities. Generally, the training process of MLLMs mainly includes two stages: the pre-training stage and the instruction tuning stage, with recent studies [63, 33, 52] primarily focusing on the second stage to enhance models’ instruction-following abilities. However, this stage gradually faces inevitable computational overhead due to the growing volume of multi-modal instruction data [36]. It is critical to explore multi-modal data selection to identify a small subset of high-quality instructions, thereby improving MLLM fine-tuning efficiency.
@@ 2.2 Instruction-based Data Selection
Although MLLMs have demonstrated remarkable performance across various tasks, data redundancy is becoming increasingly apparent as the volume of data grows exponentially, similar to what is observed with LLMs [66, 7, 12]. Previous works mainly focus on using pre-defined rules [30, 7], human feedback [37], or gradient-based approximation during training [4, 54] to select a high-quality coreset [55] while achieving competitive performance. However, these data selection methods in LLMs only aim to align the instance values of selected samples with the overall dataset, failing to effectively distinguish between similar samples or noisy data in more complex multi-modal instructions. This undoubtedly undermines the uniqueness and representativeness of the samples. For data selection in MLLMs, TIVE [36] first identifies severe redundancy in multi-modal datasets and selects valuable data at the task and instance level through gradient similarity. However, it requires extra training on downstream tasks. SELF-FILTER [53] attaches an additional evaluation model and simultaneously updates its parameters during training to select high-value samples. InstructionGPT-4 [51] selects a subset of 200 instructions for training MiniGPT4 [67], but it is unscalable for other settings. Although designed for multi-modal data selection, these methods largely follow prior approaches and overlook the complex relationships between samples, ultimately limiting the model’s generalization ability. To mitigate these limitations, we are the first to adopt a systematic perspective for multi-modal data selection by proposing three core principles: informativeness, uniqueness, and representativeness, and leveraging corresponding metrics to collaboratively assess and optimize data selection.
@@ 3 Method
As illustrated in Figure 3, our DataTailor framework consists of four primary steps: (1) The informative value captures the information density in latent space, directly reflecting informativeness to enhance MLLM generalization. (2) The unique value identifies distinct samples within the intra-cluster space, reflecting the uniqueness of sample relationships to effectively reduce redundancy. (3) The representative value captures samples that align closely with the overall dataset distribution in the inter-cluster space, ensuring representativeness and preventing compromise by noisy outliers. (4) Finally, DataTailor adaptively integrates these three values to enable collaborative multi-modal data selection. Next, we will elaborate on the details of each step.
@@ 3.1 Problem Formulation
We formulate multi-modal data selection as achieving the best performance with fewest samples by selecting a subset $S^{*}=\{s_{1},...,s_{k}\}$ with the highest value from the dataset $S$, where $s_{i}=(X_{v},X_{instruct},X_{a})$, $S^{*}\subset S$ and $k$ is the total selection proportion for dataset $S$. Therefore, the selected subset should be efficient and effective, ensuring that models trained on a limited dataset can achieve competitive performance compared to full fine-tuning. Building on our previous analysis, we systematically leverage the principles of informativeness, uniqueness, and representativeness to identify the most valuable samples.
@@ 3.2 Informative Value Estimation in Latent Space
Although several approaches have been proposed for efficient instruction-based data selection [30, 66, 47, 54], they are either infeasible at scale due to computational overhead or are limited in generalizability by pre-defined evaluation rules. To thoroughly understand the contribution of multi-modal instruction data for MLLM generalization, DataTailor directly extracts the representation of each sample within latent space to estimate its intrinsic information density, which is more accessible than previous works.
However, it is non-trivial to quantify an effective informative value to reflect the information density of samples. Cognitive Load Theory [46] asserts that both excessive and insufficient information negatively impact task generalization. Drawing upon previous research on the singular value spectrum [58, 9, 8], we propose singular value entropy (SVE) to capture the uniformity of the singular value distribution in each sample’s representation. It reflects the intrinsic information density of a sample, providing a more comprehensive measure of its informativeness by ensuring the diversity and coverage of its information dimensions. In this manner, a higher SVE value of a sample indicates that it encompasses sufficient and well-balanced information, which is crucial for robust generalization during MLLM fine-tuning. Formally, given a multi-modal instruction sample $s_{i}=(X_{v},X_{instruct},X_{a})$, we extract its unified feature matrix from the second-to-last layer $\mathbf{M_{i}}=(H_{v};H_{q})\in\mathbb{R}^{L_{i}*d}$, where $L_{i}$ is the total length of multi-modal tokens and $d$ is the feature dimension. Here, $H_{v}$ and $H_{q}$ represent the visual and instruction features, respectively. Subsequently, we perform singular value decomposition on the feature matrix in latent space $\mathbf{M_{i}}=\mathbf{U_{i}}\hat{\mathbf{\Sigma_{i}}}\mathbf{V_{i}}^{\top}$ and its corresponding diagonal singular matrix is defined as follows:
$\hat{\mathbf{\Sigma_{i}}}=\{\sigma_{0},\sigma_{1},...,\sigma_{L_{i}}\}$ (1)
where we assume $L\leq d$ [34] and all singular values $\{\sigma_{j}\}_{j=0}^{L_{i}}$ are listed in order. Building on this, we compute the entropy of normalized singular values as the informative value to assess the information density of each data:
@@
The informative value measures sample diversity and coverage across informative dimensions and thus enhances the informativeness of selected data for MLLM generalization.
3.3 Unique Value in Inter-cluster Space
To enrich the value of samples for multi-modal data selection, it is crucial to emphasize the unique values derived from their intricate relationships to reduce redundancy. The unique value estimation in the inter-cluster space consists of two steps: cross-modal domain clustering and unique value calculation. Next, we introduce each step in detail. Cross-modal Domain Clustering. Directly quantifying the unique relationships among all multi-modal instructions can be computationally expensive. To address this, we propose Cross-modal Domain Clustering to reduce the computational overhead by clustering samples within task-specific domains. Since multi-modal instruction data covers a variety of tasks, domain clustering is applied to each task category. This process generates semantically distinct clusters in each task, enhancing the diversity of the selected data. More clustering details and analyses are in Appendix B.1. Unique Value Calculation. To quantify the uniqueness of samples, we focus on identifying discriminative samples in the intra-cluster space that contribute uniquely to training. The key intuition is that discriminative samples exhibit greater distances in the cluster (as shown in Fig. 2), effectively mitigating data redundancy. Therefore, we introduce a distance coefficient to assign high unique values to these distinctive instructions based on their distance from surrounding samples, enhancing uniqueness as follows:
$V_{i}^{Uni}=\frac{1}{|C|-1}\sum_{s_{j}\in\mathbf{C},j\neq i}\|\mathbf{p_{j}}-\mathbf{p_{i}}\|_{2}\cdot V_{j}^{Inf}$ (3)
where $\|\mathbf{p_{j}}-\mathbf{p_{i}}\|_{2}$ is the Euclidean distance of two multi-modal intra-cluster instructions within latent space. In this manner, these distinctive or challenging instructions are more likely to be selected due to their enhanced unique values derived from intra-cluster relationships, thereby alleviating the issue of redundant sample selection.

### v2

@@ Abstract
Instruction tuning fine-tunes pre-trained Multi-modal Large Language Models (MLLMs) to handle real-world tasks. However, the rapid expansion of visual instruction datasets introduces data redundancy, leading to excessive computational costs. We propose a collaborative framework, DataTailor, which leverages three key principles—–informativeness, uniqueness, and representativeness–—for effective data selection. We argue that a valuable sample should be informative of the task, non-redundant, and represent the sample distribution (i.e., not an outlier). We further propose practical ways to score against each principle, which automatically adapts to a given dataset without tedious hyperparameter tuning. Comprehensive experiments on various benchmarks demonstrate that DataTailor achieves 101.3% of the performance of full-data fine-tuning with only 15% of the data, significantly reducing computational costs while maintaining superior results. This exemplifies the “Less is More” philosophy in MLLM development. The code and data is available in this URL.
00footnotetext: Corresponding author.
@@ 1 Introduction
The rapid development of Multi-modal Large Language Models (MLLMs) has made promising progress on various multi-modal tasks [68, 3, 75]. A typical MLLM is developed through two main training stages: pre-training on vast image-text pairs and fine-tuning on task-specific multi-modal instructions. Notably, the fine-tuning stage is critical for enhancing the instruction-following capabilities of MLLMs. Yet, this stage can become exceedingly time-consuming due to the large-scale but low-quality instruction data. Hence, the community is interested in fine-tuning data selection methods, such that an MLLM trained on the selected subset yields comparable or even better performance.
Figure 1: (a) The Performance v.s. Selected Data Ratio on LLaVA-mix-665k of DataTailor compared with SOTA data selection methods. (b) Metric triangle among informativeness, uniqueness, and representativeness when applying IFD [36] (data-specific methods), InsTag [44] (human-reward methods), LESS [62] (gradient-based methods), and DataTailor. (c) The corresponding MLLM performance on LLaVA-mix-665k [41] of different methods.
Existing MLLM data selection methods [12, 60, 43, 25] largely follow similar ideas from the NLP community [36, 44, 74, 55, 62]. They can be divided into three main categories: (1) Data-specific methods [36, 14, 9] leverage vast hand-designed rules to select data in specific datasets, which are not flexible and robust for diverse tasks. (2) Human-reward methods [74, 44] utilize human feedback to select data, which are both time-consuming and expensive. (3) Gradient-based methods [55, 62, 39] select samples whose gradients are similar to the average training gradient of the dataset. However, they require additional training on downstream tasks, making total computation costly.
To address these deficiencies, we build a systematic data selection method for MLLM called DataTailor. We evaluate each sample with three principles and select the most valuable samples, leading to state-of-the-art MLLM performance with a fraction of data (c.f. Fig. 1). The three principles are: (1) Informativeness: a valuable sample should be informative of the hard task, e.g., If the task is reasoning, describing the movement differences between skiing and ice skating is more informative and complex than simply describing someone skiing. In Fig. 2 where each axis (heuristically) represents an orthogonal dimension of task information, points along the diagonal carry more information about the task. (2) Uniqueness: A valuable sample should be distinct from others, offering unique insights rather than prevalent commonsense knowledge (c.f. Fig. 2 near the blue dashed region in the intra-cluster space demonstrates high uniqueness). (3) Representativeness: it should be a typical sample in the data distribution. This prevents selecting noisy outliers or mislabeled samples (c.f. Fig. 2 the clusters connected by blue lines in the inter-cluster space exhibit high representativeness for the overall dataset).
@@
We further propose a practical method to measure the value of each sample against each principle. For informativeness, we take motivations from information theory [11, 10]. For each sample, we analyze the singular value distribution of its features and use the entropy of the singular values to determine if it is informative of the task. To compute uniqueness and representativeness, we first cluster the samples based on their visual and textual features. This allows efficient calculation for uniqueness, as we can simply measure the average distances of a sample to its neighbors in the same cluster, and mark those with a large distance as unique. Then we find connected clusters and mark samples in those clusters as representative. Hence, noisy or mislabeled data from a far-away cluster can be filtered.
Moreover, as multi-modal samples exhibit varying structures and complexities across diverse tasks, we propose an adaptive weight to combine the values, which removes the need for expensive hyper-parameter tuning. We also adaptively determine the proportion of selected data for each task by using the average largest singular value of samples in the task, which empirically reflects task difficulty and correlates with training robustness. Combining these techniques, DataTailor synergizes the three principles for data selection and achieves an optimal balance between data volume and model performance (as shown Fig. 1(a) red line).
To our knowledge, we are the first to explore sample relationships between multi-modal instructions systematically. Through extensive experiments, we demonstrate that DataTailor exhibits significant effectiveness in data selection for MLLMs (with less than 5% data but achieving over 95% performance). This effectiveness is further demonstrated through quantitative metrics designed for the three principles (c.f. Fig. 1(b)), proving that DataTailor improves data selection by addressing the essential aspects of each principle (detailed analyses is in Sec. 4.4 and Appendix D.1 & D.2). In contrast, other methods lack a systematic evaluation, particularly of sample relationships, leading to weaknesses in uniqueness and representativeness and resulting in suboptimal MLLM performance (c.f. Fig. 1(c)). Remarkably, when DataTailor increases the data selection ratio, multi-modal data selection can even outperform full data fine-tuning (achieve 101.3% performance with 15% data), truly exemplifying the concept of “Less is More”. Overall, our main contributions are summarized as follows:
@@ 2.1 Multi-modal Large Language Model
With the outstanding performance of LLMs in zero-shot settings, early work combining LLMs with visual modalities has demonstrated impressive visual language comprehension abilities [20, 68, 34, 33, 35, 12, 31]. Recently, more powerful MLLMs have emerged [75, 48, 47, 67, 41, 70, 66, 22, 8], which possess perceptual abilities for visual-language tasks and excellent reasoning abilities. Generally, the training process of MLLMs primarily includes two stages: the pre-training stage and the instruction tuning stage. Recent studies [69, 40, 59, 19, 18, 52] have focused on the second stage to enhance instruction-following abilities. However, this stage is gradually facing inevitable computational overhead due to the growing volume of multi-modal data [43, 51]. It is critical to identify a small subset of high-quality instructions to improve MLLM fine-tuning efficiency.
@@ 2.2 Instruction-based Data Selection
Although MLLMs have demonstrated remarkable performance, data redundancy is becoming apparent with the rapid growth of visual instruction datasets, similar to challenges in LLMs [74, 9, 16]. Previous LLM-based works mainly focus on using pre-defined specific rules [36, 9], human feedback [44], or gradient-based approximation [4, 62] to select high-quality data. INSTRUCTMINING [9] uses 9 indicators to fit specific rules for data selection. In contrast, DataTailor automatically assesses values without hand-designed rules, offering greater robustness and flexibility. Moreover, these methods only focus on individual sample values while neglecting similar or noisy data when applied to more complex multi-modal instructions. For MLLM-based data selection, TIVE [43] identifies redundancy in multi-modal instructions and selects valuable data at the task and instance level through gradient similarity, while ICONS [61] extends this by integrating specialist influence estimation. However, they both require extra training on downstream tasks for data selection. SELF-FILTER [60] attaches evaluation models and updates its parameters during training to select high-value samples. InstructionGPT-4 [58] selects 200 instructions for MiniGPT4 [75], but it is unscalable for other settings. COINCIDE [30] roughly clusters data by conceptual representations, while overlooking the value differences between clusters. These methods largely follow prior approaches and overlook the complex relationships between samples, which limits their generalization. To mitigate these limitations, we are the first to adopt a collaborative perspective for multi-modal data selection, focusing on three core principles: informativeness, uniqueness, and representativeness.
@@ 3 Method
As illustrated in Figure 3, our DataTailor framework consists of four primary steps: (1) The informative value captures the information density in latent space, directly reflecting informativeness to enhance MLLM generalization. (2) The unique value identifies distinct samples within the intra-cluster space, reflecting the uniqueness of sample relationships to reduce redundancy effectively. (3) The representative value captures samples that align closely with the overall dataset distribution in the inter-cluster space, ensuring representativeness and preventing compromise by noisy outliers. (4) Finally, DataTailor adaptively integrates these three values to enable collaborative multi-modal data selection. Next, we will elaborate on the details of each step.
@@ 3.1 Problem Formulation
Given a visual input $X_{v}$ and a textual query $X_{q}$, the objective of the MLLM is to predict the correct answer $X_{a}$ for downstream tasks. We formulate multi-modal data selection as selecting fewest samples $S^{*}=\{s_{1},...,s_{k}\}$ from the candidate dataset $S$ for MLLM training that achieves the best inference performance, where $s_{i}=(X_{v},X_{q},X_{a})$ and $k$ is the total data selection proportion for $S$. Typically, the sample value is measured by the performance perturbations to the MLLM when removing the sample [27], but this is not practical for large-scale datasets. Therefore, we propose three core principles (i.e., informativeness, uniqueness, and representativeness) that practically identify the most valuable samples to improve MLLM’s performance on downstream tasks. More concrete instantiations of each principle are in Appendix D.2. Then, we introduce each principle.
@@ 3.2 Informative Value Estimation in Latent Space
Although several approaches have been proposed for multi-modal data selection [61, 43, 58, 60], they are either infeasible at scale due to computational overhead or are limited in generalizability by pre-defined evaluation rules. Previous studies [38, 42, 62] have shown that difficult samples contribute more to improving the performance of downstream tasks in MLLMs. Thus, DataTailor directly estimates samples’ essential difficulty for multi-modal data selection.
However, it is non-trivial to quantify sample difficulty for multi-modal instructions due to varying information representations in different modalities. Drawing upon previous spectral analysis [65, 11, 10], we perform singular value decomposition (SVD) on the unified feature matrix at the token level for image and text modalities to reflect the difficulty of samples, as it consists of the variations across different tokens in latent space. Formally, given a multi-modal sample $s_{i}$, we obtain its unified feature matrix $\mathbf{M_{i}}\in\mathbb{R}^{L_{i}*d}$ from the penultimate layer, where $L_{i}$ is the token length and $d$ is the feature dimension. We adopt this layer as the lower layers struggle to capture complex features, while the final layer suffers biases of pre-training datasets [42]. Besides, the penultimate layer retains earlier layer information, providing richer features. Subsequently, we perform SVD on the matrix in latent space $\mathbf{M_{i}}=\mathbf{U_{i}}\hat{\mathbf{\Sigma_{i}}}\mathbf{V_{i}}^{\top}$ and its corresponding diagonal singular matrix is defined as follows:
$\hat{\mathbf{\Sigma_{i}}}=\{\sigma_{1},...,\sigma_{L_{i}}\}$ (1)
where we assume $L_{i}\leq d$ [41] and all singular values $\{\sigma_{j}\}_{j=1}^{L_{i}}$ are listed in order. Building on this, we compute the entropy of normalized singular values as the informative value to assess the information density of each data:
@@
Intuitively, simple samples contain redundant information in images or text, and only part of information is necessary to answer the query (see Appendix D.2 for examples). As a result, the columns of their feature matrices exhibit strong linear dependence, which drive down the smaller singular values, causing the top ones to be much bigger by comparison, leading lower singular value entropy. In contrast, more challenging samples are richer in information, with feature matrices that are closer to full rank, resulting in more uniform singular values and higher entropy. In this way, SVE can as a practical approximation to measure the difficulty of samples, enabling the selection of hard samples to enhance the performance of MLLM in downstream tasks.
3.3 Unique Value in Intra-cluster Space
To enrich the value of samples for multi-modal data selection, it is crucial to emphasize the unique values of samples to select unique samples from similar clusters. The unique value estimation in the intra-cluster space consists of two steps: cross-modal domain clustering and unique value calculation. Next, we introduce each step in detail.
Cross-modal Domain Clustering. To efficiently distinguish unique samples, we propose an optimized Cross-modal Domain Clustering that aggregate sufficiently similar samples into cluster in each task. This process begins with quantifying the $\ell_{2}$-norm distance between samples for cluster variance computation. It then utilizes the Ward criterion [46] to merge clusters with the minimum increase in sum of squared errors (SSE) progressively, computed as:
$\Delta\text{SSE}=\frac{n_{A}\cdot n_{B}}{n_{A}+n_{B}}\cdot\|\bm{\mu}_{A}-\bm{\mu}_{B}\|_{2}$ (3)
where $n_{A},n_{B}$ denote cluster cardinalities and $\bm{\mu}_{A},\bm{\mu}_{B}$ represent cluster centroids. The clustering terminates when the $\Delta\text{SSE}$ exceeds the threshold $\lambda\Delta\text{SSE}_{\text{max}}$, where $\Delta\text{SSE}_{\text{max}}$ denotes the maximum increase in SSE. Here, $\lambda$ controls the number of clusters and is set to 0.1 by default to match the total data selection proportion for capturing sample relationships. Moreover, we use parallel computation and memory optimization for efficiency. More implementation details and analyses are shown in Appendix B.3.
Unique Value Calculation. To quantify the uniqueness of samples, we focus on identifying discriminative samples in the intra-cluster space that contribute uniquely to training. The key intuition is that unique samples exhibit larger distance from similar samples (as shown in Fig. 2) to contain more information in the cluster for generalization. Therefore, we introduce a distance coefficient to assign high unique values to these distinctive instructions based on their distance from surrounding samples as follows:
$V_{i}^{Uni}=\sum_{s_{j}\in\mathbf{C},j\neq i}\|\mathbf{p_{j}}-\mathbf{p_{i}}\|_{2}\cdot\frac{V_{j}^{Inf}}{\sum_{k\in\mathbf{C}}V_{k}^{Inf}}$ (4)
where $\|\mathbf{p_{j}}-\mathbf{p_{i}}\|_{2}$ is the Euclidean distance of two samples $s_{i},s_{j}$ in the cluster $\mathbf{C}$ in latent space. We assign higher reference weights to samples with greater informative value when computing their unique and representative values, reflecting their stronger influence on other samples. In this manner, these challenging instructions are more likely to be selected due to their enhanced unique values.